meatqueen:

I'm solving equations using square roots. g2 = 25????

ramen:

You can't give out direct answers batman but basically, what is the square root of 25? What number times itself=25?

ramen:

This is for the value of g.

Imagine:

What I was going to say..

Imagine:

\[\sqrt{25}\]

Imagine:

\[5 \times 5=?\]

jhonyy9:

g^2 = 25 \[g_1,_2 = +/- \sqrt{25} = ?\]

jhonyy9:

@Imagine

Imagine:

True!! I have to pay more attention!

jhonyy9:

not just + 5 satisfy this

Imagine:

It could be \[5~or -5\]

Imagine:

Thanks Jhonny!

jhonyy9:

not or and

jhonyy9:

every two are right,correct answer

ramen:

Anytime you use a square root, it is + or -.

jhonyy9:

not or + and -

ramen:

Yep that's what I was tryna say. Thanks

ramen:

Because -5*-5=25 while 5*5=25.

meatqueen:

Thanks. I guess

Imagine:

Very True! Thanks Jhonny for pointing that out!

ramen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @meatqueen Thanks. I guess \(\color{#0cbb34}{\text{End of Quote}}\) np